OpenStudy (anonymous):

Write an equation of a parabola that opens upward, has a vertex at the origin, and a focus at (0,2)

OpenStudy (apoorvk):

The standard equation of a parabola with vertex at origin and it's axis vertical is: \[x^2 = 4ay\] and the focus of this lies at (0,a) So, find out 'a', since you have the coordinates of the focus.

OpenStudy (anonymous):

in old notation parabola looks like \[4p(y-k)=(x-h)^2\] since your vertex is at the origin, this becomes \[4py=x^2\] and now all you need is \(4p\) since \(p=2\) you have \(4p=8\) and your parabola is \[8y=x^2\]